Alicia is thinking of a number. She subtracts the square of this number from six times the number and then subtracts eight from the result.
a) use x to write the expression that describes the information
b) factorise the expression in part a)
c) what could this help us to find out?

Answer by ankor@dixie-net.com(22740)

She subtracts the square of this number from six times the number and then subtracts eight from the result.
a) use x to write the expression that describes the information
-x^2 + 6x - 8 = 0
b) factorise the expression in part a)
Easier to factor if you the coefficient of x^2 is one, mult by -1
x^2 - 6x + 8 = 0
(x-4)(x-2) = 0
two solutions
x = 4
x = 2
c) what could this help us to find out? Not sure what this means here.